Shell environment for furoligos.tex

Thomas D. Schneider^{∗}

version = 1.00 of shell.tex 2006 Oct 26

∗National Cancer Institute at Frederick, Center for Cancer Research Nanobiology Program, P. O. Box B, Frederick, MD 21702-1201. (301) 846-5581 (-5532 for messages), fax: (301) 846-5598, email: toms@ncifcrf.gov. http://www.lecb.ncifcrf.gov/˜ toms/

Name  Begin  End  Length Sequence
(base)  (base)  (bases) (5′to 3′)
yoeA  2066596 2066638 43 aacataaatgaaaataattatcattacagtaatcatttgtact
fepD-entS 621416 621458  43 aacttccatgataatgaaattaattatcgttatcgatcttatt
gpmA 786868 786838 31 caatataatgagaattattatcattaaaaga
ryhB  3579039 3579069 31 cgagacaataataatcattctcattcgcact
fhuA 167418 167454 37 ttatcttatctttataataatcattctcgtttacgtt
nohA  1634639 1634609 31 tatgtagatgataatcattatcactttacgg
oppA  1298952 1298988 37 ttctttattggtaatgagaataattaacaattaaaga
gspC  3453690 3453720 31 attatttttaatgtgaattatttccatacaa
yhaU  3272849 3272882 34 ggtaacaataaaaataatcagtaatattaaatag
yahA 331880 331916 37 cacagatatattaatgataatcactatcaccatatcg
fadD  1886636 1886666 31 taccactatgataatcaatatcatatgggtt
ygaC  2798602 2798632 31 tccagggtaatttcgaccactatttgctata
exbB  3150057 3150103 47 acgacgtcgaattgatagtcgttctcattactatttgcatactgccg
exbD  3149440 3149486 47 aatcgctgccacgaggccgattgccgttgctaacagagcttctgcga
fhuF  4603768 4603815 48 caaatgataaaaattatcatatgatattggttatcattatcaatgaaa

Table 1: Sequences of natural Fur sites used for oligo constructions.

The sequence is from E. coli K12, GenBank Accession NC 000913. See Materials and Methods for how these sequences were used. Sequence walker analysis of the oligos is available at http://www.ccrnp.ncifcrf.gov/∼toms/papers/fur/.